OpenStudy (anonymous):

I need help with the equation for this... You have a batch of squared tiles. If you make with them a square of 'x' tiles per side, 27 tiles remain, and if you take 'x+1' tiles per side, 40 are needed. Find how many tiles are on the batch.

OpenStudy (anonymous):

(x+1)^2 = 40 and plug the answer to that into x^2+27 to get the batch size

OpenStudy (anonymous):

so...the equation plugged it's \[x^2-38=x^2+27\] ?

OpenStudy (anonymous):

It's not going to be pretty because 40 is not a square Solve for X in the first equation and get 2sqrt(10) - 1 then square it and add 27 and get 68 - 4sqrt(10)